OpenStudy (kayleigh98):

Write the equation for the line using the given information. Write the equation in slope-intercept form, y = mx + b. Write the equation of the line with slope = 5 and passing through the point (−4, −7).

OpenStudy (lynfran):

you already have all the info to solve this... slope=5 so... m=5 points (x, y) so,.. x = -4 and y=-7 now sub those values into the given slope - intercept formula ... then solve for "b" which is the y intercept

OpenStudy (lynfran):

what equation did u get ..?

OpenStudy (kayleigh98):

So i take those and substitute them into the equation for b?

OpenStudy (lynfran):

yes like this y=mx+b -7=5(-4)+b now solve for "b"

OpenStudy (kayleigh98):

7/20?

OpenStudy (lynfran):

no... first we multiply the parentheses ..we get -7=-20+b then we add 20 to both sides of the equation ...we get -7+20=-20+20+b simplify we get 20-7=b ...whats b=?

OpenStudy (kayleigh98):

13?

OpenStudy (kayleigh98):

great i got it thank you!

OpenStudy (lynfran):

yes and thats the y intercept now we write the equation using the slope and y intercept y=mx+b ... we just sub the value of "m" and "b"